OpenStudy (anonymous):

Help! just double checking to see if i am right. What is the P(Red l Heart) when drawing one card?

OpenStudy (anonymous):

Perv^ Since a Heart is a red card the numerator can be simply written as P(Red | Heart) = P(Heart) / P(Heart) = [13/52] / 13/52 = 13/52 * 52 / 13 = 1

OpenStudy (anonymous):

there is 13 red hearts in a card deck and there is 52 cards all together thats how I got 13 and 52

OpenStudy (anonymous):

gtbishawt this is a place for homework not a place for getting a girl friend ps she is smooking hot i just hope i see her in real life i will mess her up after we have sex we will have it agian non stop bam she will be sucking my wingspan in front of your face

OpenStudy (anonymous):

Eew... and thanks! Cpt.Sparz!!:) Can you hlep with something else?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

It's all good they got no life xD

OpenStudy (anonymous):

what girl you know you want mee

OpenStudy (anonymous):

.... you are disgusting creatures...

OpenStudy (anonymous):

:D thanks! What is the P(Jack ∩ Club) when drawing one card?

OpenStudy (anonymous):

ya you know i am i cant wait to see ya

OpenStudy (anonymous):

No problem ! P(Jack | Face Card) = P( Jack and Face Card) / P(Face Card) A Jack is a face card so the numerator of the above fraction can be written simply as P(Jack | Face Card) = P(Jack) / P(Face Card) P(Jack) = 4/52 P(Face Card) = 12/52 So P(Jack | Face Card) = (4/52) / (12 / 52) = 4/52 * 52/12 = 4/12 P(Jack | Face Card) = 1/3

OpenStudy (anonymous):

Makes so much more sense! lol Thank you so much! lol mind if i ask you two more things??

OpenStudy (anonymous):

sup goergus

OpenStudy (anonymous):

na ask away

OpenStudy (anonymous):

lol no problem and sure :D

OpenStudy (anonymous):

dude she was talking to mee

OpenStudy (anonymous):

lol alright mate

OpenStudy (anonymous):

dude what grade is you in

OpenStudy (anonymous):

:D yay! thanks - Cpt.Sparz! Choose an employed person at random. Let A be the event that the person is female and B be the event that the person holds a managerial position. Data from the U.S. Department of Labor suggests that P(S) = 0.47 and P(BlA) = 0.34. a) Explain what P(A) = 0.47 means in context b) Explain what P(BlA) = 0.34 means in context

OpenStudy (anonymous):

I'm a junior XD

OpenStudy (anonymous):

To be honest idk this one... I:

OpenStudy (anonymous):

This will take sometime, but sometimes i get these right.

OpenStudy (anonymous):

How much time can i get to try to solve this?

OpenStudy (anonymous):

Nice Asking Alexandria shirt btw! :D only junior i know with good taste in music.

OpenStudy (anonymous):

lol that's fine. i appreciate the attempt in helping though:)

OpenStudy (anonymous):

As long as you need! FlvsGirl

OpenStudy (anonymous):

Yea, this question is like not working out with me

OpenStudy (anonymous):

haha its fine.. i appreciate the attempt though :)

OpenStudy (anonymous):

ok? what else do you need?

OpenStudy (anonymous):

lol Ty gtb

OpenStudy (anonymous):

No problem! :D

OpenStudy (anonymous):

still trying to figure out the problem sorry it's taking so long

OpenStudy (anonymous):

@Cpt.Sparz Same here >.<

OpenStudy (anonymous):

lol yea it's kind of hard

OpenStudy (anonymous):

Haha, yea my brain is dead on this one.....Sorry to the one who asked it, I tried but I cant come to a right answer and explaination :/

OpenStudy (anonymous):

Yea same it's irritating when I can't help at all

OpenStudy (anonymous):

@Cpt.Sparz So true, hey are you able to acceptp messages? We can collaberate on this problem, if not cool

OpenStudy (anonymous):

That's completely fine you guys. :) Don't beat yourselves up over it, i just appreciate your attempt at helping!:)

OpenStudy (anonymous):

yea I can accept messages

OpenStudy (anonymous):

@gtb0329 Thanks....Yea, im puzzled.....

OpenStudy (anonymous):

@Cpt.Sparz Cool, yea I am like really troubled with the problem. :/

OpenStudy (anonymous):

lol yeah, its definitely a frustrating problem

OpenStudy (anonymous):

Yea same here

OpenStudy (anonymous):

@gtb0329 I fanned you, and thanks for being cool over it all....

OpenStudy (anonymous):

haha thanks! XD and it's no problem at all.. just glad you've tried :)

OpenStudy (anonymous):

Math isnt my strongest subject......though I push myself to learn it and know it

OpenStudy (anonymous):

Yea math i'm good at but not the best

OpenStudy (anonymous):

I'm really good, but only if i am taught properly

OpenStudy (anonymous):

Dont worry, I will not rest till I get it....Never Shout Never..... I love the guy/group Never Shout Never

OpenStudy (anonymous):

nice lol.. they remind me of the rocket summer

OpenStudy (anonymous):

I think I did it, I dont know P(Red | Heart) = P(Red card and Heart) / P(Heart) Since a Heart is a red card the numerator can be simply written as P(Red | Heart) = P(Heart) / P(Heart) = [13/52] / 13/52 = 13/52 * 52 / 13 = 1 P(Red | Heart) = 1 This makes sense, since if you knew you that the card you picked was a heart (which is a red card) you would be absolutely certain that you picked a red card.

OpenStudy (anonymous):

It took me 4everrrrrr, but I had to keep fixxing it

OpenStudy (anonymous):

i meant romance on a rocketship, but they work too.. and yeah, Cpt.Sparz already gave me that :P But it is right!:D

OpenStudy (anonymous):

@Cpt.Sparz How did i do? I think I did okay, but not great

OpenStudy (anonymous):

Same thing really lol good job :D

OpenStudy (anonymous):

Thank you, lol....My brain can not rest, though not for long for I have to do spanish, and science

OpenStudy (anonymous):

hahaha no problem I can help you in spanish if you like

OpenStudy (anonymous):

haha yeah, i got a little feeling of Déjà vu with that... Ugh, i was so glad to get out of that class last year! it frustrated me so much. I can attempt to help you since i've already finished :D

OpenStudy (anonymous):

@Cpt.Sparz Thanks sooo much...where though? I dont think there is a section for it

OpenStudy (anonymous):

There is lol I've taken spanish 1 and 2...

OpenStudy (anonymous):

yeah, i quit after that, i was not in the mood for a spanish 3

OpenStudy (anonymous):

Meet me in the Spanish section, I will go type the questions...Spanish 1, sadly.